OpenStudy (anonymous):

Simplify i41 i −1 −i 1

OpenStudy (anonymous):

i^41

OpenStudy (anonymous):

i assume this is \[i^{41}\]

OpenStudy (anonymous):

yess

OpenStudy (anonymous):

\[i^0=1\\ i^1=i\\ i^2=-1\\ i^3=-i\\ i^4=1\] and the pattern continues

OpenStudy (anonymous):

take the integer remainder when you divide \(41\) by \(4\) you get \(1\) since \(4\) goes in to \(40\) \(10\) times with a remainder of \(1\) therefore \(i^{41}=i^1=1\)

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

\[i^{41}=i^1=i\] is the correct answer

OpenStudy (anonymous):

-1?